OpenStudy (anonymous):

Find the sum of a finite geometric sequence from n = 1 to n = 6, using the expression −2(5)^(n − 1)

hartnn (hartnn):

−2(5)^(n − 1) compare it with \(\Large a_1 r^{n-1}\) you will get a1 and r

OpenStudy (anonymous):

a1=-2 and r=5

hartnn (hartnn):

then use the sum formula \(\Large S_n =a_1 \dfrac{r^n-1}{r-1} \)

OpenStudy (anonymous):

Sn=-2(5^(n)-1/5-1)?

OpenStudy (anonymous):

sorry my equation button isn't working at the moment

hartnn (hartnn):

no problem and thats correct simplify that

hartnn (hartnn):

plug in n=6 because u need sum of 6 terms

OpenStudy (anonymous):

so Sn=-2(5^(6)-1/5-1?

hartnn (hartnn):

-2 [(5^6 -1)/ (5-1)]

OpenStudy (anonymous):

yes that

hartnn (hartnn):

simplify it

OpenStudy (anonymous):

-2*(3906)

OpenStudy (anonymous):

-7812

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

yay thank you so much!